alexisHales:

The area of a trapezoid is calculated using the formula below, where A is the area of the trapezoid, b1 and b2 are the bases of the trapezoid, and h is the height of the trapezoid. Rewrite the formula to find the base b2. a. b2=2A-h-b1 b. b2=(2A/h)-b2 c. b2=2A*h-b1 d. b2=2((A/h)-b1)

dude:

Welcome to QuestionCove! Could you make sure to include the equation given in the question?

alexisHales:

oh sorry A=(b1+b2/2*h

alexisHales:

A=(b1+b2/2)*h

dude:

\(A=(\dfrac{b1+b2}{2})*h\) Does this look right?

alexisHales:

yea

alexisHales:

._.

dude:

Okay, so we basically want to move all the other variables to the other side We do this by performing the opposite operation. Let's start with h, how would you move h to the other side?

alexisHales:

division

dude:

Right, we divide h on both sides We get, \(\dfrac{A}{h}=\dfrac{b1+b2}{2}\) Now we can get rid of the 2, do you know how to do that?

alexisHales:

multiplication leaving you with only b1+b2 on the right sight. am right?

dude:

Whoops, my bad \(\dfrac{2A}{h}={b1+b2}\)

alexisHales:

is that all?

dude:

Now we can isolate b2 Thats out last step, do you know how to do that?

alexisHales:

subtract b1?

dude:

Yeah

alexisHales:

so b1+(2A/3)=b2 ???

alexisHales:

wait i mean subtract sign lol

alexisHales:

b1-(2A/3)=b2 b2=(2A/3)-b1

dude:

Oh 2 was written the other way \(2\dfrac Ah-b1\) In this case they decided as \(2((\dfrac Ah)-b1)\)

alexisHales:

oh ok